# Interpretable classifiers for tabular data via feature selection and discretization

### v2

@@ Abstract
We introduce a method for computing immediately human interpretable yet accurate classifiers from tabular data. The classifiers obtained are short Boolean formulas, computed via first discretizing the original data and then using feature selection coupled with a very fast algorithm for producing the best possible Boolean classifier for the setting. We demonstrate the approach via 13 experiments, obtaining results with accuracies comparable to ones obtained via random forests, XGBoost, and existing results for the same datasets in the literature. In most cases, the accuracy of our method is in fact similar to that of the reference methods, even though the main objective of our study is the immediate interpretability of our classifiers. We also prove a new result on the probability that the classifier we obtain from real-life data corresponds to the ideally best classifier with respect to the background distribution the data comes from.
@@ 1 Introduction
The classifiers we produce are given in the form of short Boolean DNF-formulas. One of the key issues that makes our approach possible is the surprising power of successful feature selection. Indeed, in our formulas, we use very small numbers of attributes, making them short and thus interpretable. Already in [12], it was observed that by using a single attribute, one could get, for many of the 16 datasets studied in that paper, an accuracy not drastically different from the then state-of-the-art decision trees. We apply this “simplicity first” approach in our method, together with a very rough discretization of numerical attributes to Boolean form. While we focus on interpretability, we obtain also relatively accurate classifiers. Indeed, our classifiers perform surprisingly well in comparison to widely recognized methods despite the fact that generally, there exists an obvious trade-off between accuracy and interpretability.
@@ 1.1 Overview of our method and contributions
As already indicated, we test our method on 13 tabular datasets: seven binary classification tasks from the UCI machine learning repository; a high-dimensional binary classification task with biological data; and five benchmark sets from [10]. See Section 4 for a detailed list. In each test, we compare the accuracy of our classifier to a result obtained by state-of-the art classifiers. Now, taking into account that tabular data is still a challenge to deep learning methods [10, 3], we use both XGBoost [6] and Sklearn’s implementation of random forests [4] as reference methods. These models are widely recognized for their performance on tabular data [22, 10]. For the UCI datasets we additionally compare to the method of [14] based on the length of formulas. This method is computationally too inefficient to use on the other datasets we consider. In addition to the reference methods, in relation to 10 of the experiments, we also report results found in the literature for the used datasets.
For 12 datasets we use standard ten-fold cross-validation and report the average accuracy as well as the standard deviation over the ten splits. For the high-dimensional dataset Colon with less than 100 data points we instead use leave-one-out cross-validation.
@@
In addition to local interpretability, most of the formulas produced in the experiments are in fact so simple that they can be readily globally interpreted, meaning that their behaviour on any input is clear simply by the form of the formula.
@@
In relation to accuracy, the experiments are summarized in Figures 1 and 2. In Table 4.2 we report the average number of attributes (features) used by the final formulas and the dimensions of the datasets.
In general, our method has similar accuracy to the reference methods. For the UCI datasets reported in Figure 1, our method obtains accuracies on par with and sometimes better than the state-of-the-art reference methods. For example, on the Hepatitis dataset we obtain an average accuracy of 80.7 percent compared to the 78.2 percent of random forests, 79 percent of XGBoost and 79.4 percent of the formula-size method of [14]. The five benchmark datasets from [10] reported in Figure 2 were the most difficult for our method, but even for these we obtained some surprisingly accurate and reasonably interpretable formulas. For example we obtain a formula of the form
@@
Concerning our experiments, we stress once more that the main advantage of out method is interpretability. As Table 4.2 indicates, most of our classifiers use very small numbers of features, thus being interpretable. Some classifiers obtained in the tests are longer, but still within the bounds of local (while not necessarily global) interpretability.
Finally, a key insight behind our method is the idea of computing the ideal classifiers (i.e., the above mentioned best possible Boolean classifiers) and the fact that this can be done fast using the $\mathcal{O}(|W||\tau||T|)$ algorithm (where $|T|\leq\min(|W|,2^{|\tau|})$) when $\tau$ is small. Since small $\tau$ is often sufficient—which is another key insight in the method—the approach indeed works quite accurately and fast. The ideal classifiers being central to the approach, we call the method the ideal classifier method. A possible alternative for this is would be the ideal DNF method.
@@ 3 Feature selection and overfitting
    Suppose one procedure selected the set $F=\{p_{i_{1}},\dots,p_{i_{\ell}}\}$ of features. We use Algorithm 1 to compute the empirical ideal classifier for $q$ on the data $W_{\mathrm{train}}$ using the attributes in $F$. Note that this is the best possible classifier for $q$ over $F$, and given in DNF-form. We do this step for all three sets of features given by the procedures. We select the formula with the highest validation accuracy $r$ on the data $W_{\mathrm{val}}$ and record the tuple $(\ell,F,r)$ for future steps.
@@ 4.1 Experimental setup
20202020404040406060606080808080100100100100BankMarketingBreastCancerCongressionalVotingGermanCreditHeartDiseaseHepatitisStudentDropout96.696.696.696.696.396.396.396.378.078.078.078.081.681.681.681.687.287.287.287.288.988.988.988.995.595.595.595.596.396.396.396.370.170.170.170.173.673.673.673.679.479.479.479.480.480.480.480.489.089.089.089.097.197.197.197.195.495.495.495.474.774.774.774.780.880.880.880.879.079.079.079.087.487.487.487.489.189.189.189.197.497.497.497.496.196.196.196.174.374.374.374.381.881.881.881.878.278.278.278.286.786.786.786.788.888.888.888.895.995.995.995.996.396.396.396.370.870.870.870.881.281.281.281.280.780.780.780.781.181.181.181.1%Our methodRandom forestXGBoostFormula sizeUCI repo
Figure 1: The average test accuracies and standard deviations obtained for the UCI datasets with our method, random forests, XGBoost and the formula-size method. Where available, we include also accuracies reported as Baseline Model Performance in the UCI repository, although these are not directly comparable due to different technical particularities in the experiments.
We compared our method empirically to random forests and XGBoost. These two comparison methods are very commonly used and state-of-the-art for tabular data. We tested all three methods on $13$ tabular datasets. The datasets can be categorized as follows.
@@
    $5$ tabular-data benchmarks out of $7$ binary classification benchmarks that were presented in the paper [10]. The five benchmarks that we picked (out of the seven) were the ones without missing values. These datasets were also originally from the UCI machine learning repository.
@@
As mentioned in Section 3, our method utilizes readily available feature selection methods. We used Scikit-learn’s SelectKBest, which returns $k$ features that have the highest scores based on a given univariate statistical test. The system SelectKBest supports three methods for calculating the scores for the features: F-test, mutual information, and $\chi^{2}$-test. When generating formulas, we tested all three of these methods and selected the best in terms of test accuracy. We emphasize that feature selection was performed after Booleanizing the data.
@@ 4.2 Results
20202020404040406060606080808080100100100100COMPASCovertypeElectricityEyeMovementRoadSafetyColon84848484787878789292929289898989656565657777777772.672.672.672.667.667.667.667.697.897.897.897.893.393.393.393.367.267.267.267.283.283.283.283.283.983.983.983.967.567.567.567.597.697.697.697.688.188.188.188.165.065.065.065.080.080.080.080.080.680.680.680.665.865.865.865.874.774.774.774.772.172.172.172.155.355.355.355.373.173.173.173.1%Our methodRandom forestXGBoostLiterature
Figure 2: The average test accuracies obtained on the benchmark and biological datasets with our method, random forests and XGBoost. For the benchmark datasets also standard deviations are reported. The dimensions of these datasets are too large for the formula-size method. When available, we have also included accuracies reported in literature, though these are not directly comparable due to different technical particularities in the experiments.
Figures 1 and 2 contain the average test accuracies and standard deviations that were obtained by using our new method, random forests and XGBoost using ten-fold cross-validation. Figure 1 also contains results obtained with the formula-size method of [14]. Both figures also contain accuracies that we found in the literature for these datasets. For Figure 1, we include the accuracies that were reported for these datasets in the UCI machine learning repository. For Figure 2, for the case of Colon, we used the result reported in [23], while for remaining datasets we reported the accuries given in the full-version of [10].
@@
In both Figures 1 and 2 we see that our method produces classifiers that mostly have accuracies similar to the ones obtained by the alternative methods. In particular, on medical data such as BreastCancer and Hepatitis, our method compares well to the reference methods. Perhaps the biggest gaps are found in the cases of CoverType and Electricity, two of the benchmark datasets. However, we emphasize that in [10], it was mentioned that all of the benchmark datasets were specifically selected to be “not too easy”. Hence, one should not perhaps expect that “simple” classifiers with focus on interpretability will perform too well with these datasets.
@@
We list the average runtime of our method on a single split of the data in the Appendix 6.3. For six of the datasets, the average runtime was less than 30 seconds. Of the remaining datasets, five had runtimes of less than 12 minutes. Even for the two largest datasets, Covertype and RoadSafety, the runtime is still less than an hour and, notably, systematically less than that of the hyperparameter optimizations of random forests and XGBoost. The experiments on our method, random forests and XBGoost were run on a laptop.
@@
While in Figures 1 and 2 we report the averages and standard deviations of test accuracies obtained, in Table 4.2 we report the average of the smallest number of features sufficient to reach the reported accuracy. Table 4.2 also gives the number of rows and attributes in the Booleanized datasets as well as the original datasets. We see from Table 4.2 that for eight out of 13 datasets, the best accuracy was reached already with on average less than four features, leading to very short formulas.
@@
Next we discuss the results obtained using the formula-size method of [14]. By the size of a propositional formula they mean the number of proposition symbols, conjunctions, disjunctions and negations in the formula. The results are given as part of Figure 1.
@@
We see that for most datasets in Figure 1, the accuracies obtained by the formula-size method are similar to our method. A meaningful difference can be seen in the HeartDisease dataset, where both the accuracy of 73.6 percent and standard deviation of 12.2 percent are significantly behind other methods. The formulas obtained are short, with formula size mostly less than 10.
@@
The downside of the formula-size method is computational resources. As described in the previous subsection, we used a Linux cluster featuring Intel Xeon 2.40 GHz CPUs with 8 CPU cores per run, employing a timeout of 72 hours per instance and a memory limit of 64 GB. For all but one dataset we ran the formula-size method on, the computation took more than 24 hours per split of the data with this high efficiency setup. Even for the easiest one, BreastCancer, the computation time was in the order of 12 hours per split. Furthermore, for the benchmark datasets and the high-dimensional biological dataset Colon, the formula-size method is unfeasible already for very small formula sizes.
@@
{tblr}
—l—c—c—c—c— Data set Selected features Boolean features Original features Data points
@@ 4.3 Examples of obtained formulas
In the Appendix 6.4, there are further examples of formulas obtained as outputs of our method. For each dataset, we list the formula obtained for the first 90/10 split of our ten-fold cross-validation. While some of these formulas are quite long, they should still be rather fast to use for local explanations explicating why a particular input was classified in a particular way (cf. the Introduction for details on local explanations). For these experiments, we used the same maximum number 10 of features and the same limit of 1 percentage point to choose an accurate enough formula. These are hyperparameters that one could optimize to obtain a better balance of interpretability and accuracy on any specific dataset.

### v3

@@ Abstract
We introduce a method for computing immediately human interpretable yet accurate classifiers from tabular data. The classifiers obtained are short Boolean formulas, computed via first discretizing the original data and then using feature selection coupled with a very fast algorithm for producing the best possible Boolean classifier for the setting. We demonstrate the approach via 12 experiments, obtaining results with accuracies comparable to ones obtained via random forests, XGBoost, and existing results for the same datasets in the literature. In most cases, the accuracy of our method is in fact similar to that of the reference methods, even though the main objective of our study is the immediate interpretability of our classifiers. We also prove a new result on the probability that the classifier we obtain from real-life data corresponds to the ideally best classifier with respect to the background distribution the data comes from.
@@ 1 Introduction
The classifiers we produce are given in the form of short Boolean DNF-formulas which can naturally be conceived as simply Boolean concepts. One of the key issues that makes our approach possible is the surprising power of successful feature selection. Indeed, in our formulas, we use very small numbers of attributes, making them short and thus interpretable. Already in [12], it was observed that by using a single attribute, one could get, for many of the 16 datasets studied in that paper, an accuracy not drastically different from the then state-of-the-art decision trees. We apply this “simplicity first” approach in our method, together with a very rough discretization of numerical attributes to Boolean form. While we focus on interpretability, we obtain also relatively accurate classifiers. Indeed, our classifiers perform surprisingly well in comparison to widely recognized methods despite the fact that generally, there exists an obvious trade-off between accuracy and interpretability.
@@ 1.1 Overview of our method and contributions
As already indicated, we test our method on 12 tabular datasets: seven binary classification tasks from the UCI machine learning repository; a high-dimensional binary classification task with biological data; and four benchmark sets from [10]. See Section 4 for a detailed list. In each test, we compare the accuracy of our classifier to a result obtained by state-of-the art classifiers. Now, taking into account that tabular data is still a challenge to deep learning methods [10, 3], we use both XGBoost [6] and Sklearn’s implementation of random forests [4] as reference methods. These models are widely recognized for their performance on tabular data [22, 10]. For the UCI datasets we additionally compare to the method of [14] based on the length of formulas. This method is computationally too inefficient to use on the other datasets we consider. In addition to the reference methods, in relation to 10 of the experiments, we also report results found in the literature for the used datasets.
For 11 datasets we use standard ten-fold cross-validation and report the average accuracy as well as the standard deviation over the ten splits. For the high-dimensional dataset Colon with less than 100 data points we instead use leave-one-out cross-validation.
@@
In addition to local interpretability, most of the formulas produced in the experiments are in fact so simple that they can be readily globally interpreted, meaning that their behaviour on any input is clear simply by the form of the formula. Globally interpreting a short DNF-formula involves looking at each of the few conjunctions separately and interpreting the meaning of their particular combination of attributes. The behaviour of the entire formula is then given by the fact that it accepts the cases given by the conjunctions and rejects everything else.
@@
In relation to accuracy, the experiments are summarized in Figure 1. In Table 1 we report the average number of attributes (features) used by the final formulas and the dimensions of the datasets.
In general, our method has similar accuracy to the reference methods. For the UCI datasets reported on the left in Figure 1, our method obtains accuracies on par with and sometimes better than the state-of-the-art reference methods. For example, on the Hepatitis dataset we obtain an average accuracy of 80.7 percent compared to the 78.2 percent of random forests, 79 percent of XGBoost and 79.4 percent of the formula-size method of [14]. The four benchmark datasets from [10] reported on the right in Figure 1 were the most difficult for our method, but even for these we obtained some surprisingly accurate and reasonably interpretable formulas. For example we obtain a formula of the form
@@
Concerning our experiments, we stress once more that the main advantage of our method is interpretability. As Table 1 indicates, most of our classifiers use very small numbers of features, thus being interpretable. Some classifiers obtained in the tests are longer, but still within the bounds of local (while not necessarily global) interpretability.
Finally, a key insight behind our method is the idea of computing the ideal classifiers (i.e., the above mentioned best possible Boolean classifiers) and the fact that this can be done fast using the $\mathcal{O}(|W||\tau||T|)$ algorithm (where $|T|\leq\min(|W|,2^{|\tau|})$) when $\tau$ is small. Since small $\tau$ is often sufficient—which is another key insight in the method—the approach indeed works quite accurately and fast. The ideal classifiers being central to the approach, we call the method the ideal classifier method. A possible alternative for this is would be the ideal DNF-method.
@@ 3 Feature selection and overfitting
    Suppose one procedure selected the set $F=\{p_{i_{1}},\dots,p_{i_{\ell}}\}$ of features. We use Algorithm 1 to compute the empirical ideal classifier for $q$ on the data $W_{\mathrm{train}}$ using the attributes in $F$. Note that this is the best possible classifier for $q$ over $F$, given in DNF-form. We do this step for all three sets of features given by the procedures. We select the formula with the highest validation accuracy $r$ on the data $W_{\mathrm{val}}$ and record the tuple $(\ell,F,r)$ for future steps.
@@ 4.1 Experimental setup
20202020404040406060606080808080100100100100BankMarketingBreastCancerCongressionalVotingGermanCreditHeartDiseaseHepatitisStudentDropout96.696.696.696.696.396.396.396.378.078.078.078.081.681.681.681.687.287.287.287.288.988.988.988.995.595.595.595.596.396.396.396.370.170.170.170.173.673.673.673.679.479.479.479.480.480.480.480.489.089.089.089.097.197.197.197.195.495.495.495.474.774.774.774.780.880.880.880.879.079.079.079.087.487.487.487.489.189.189.189.197.497.497.497.496.196.196.196.174.374.374.374.381.881.881.881.878.278.278.278.286.786.786.786.788.888.888.888.895.995.995.995.996.396.396.396.370.870.870.870.881.281.281.281.280.780.780.780.781.181.181.181.1%
20202020404040406060606080808080100100100100CovertypeElectricityEyeMovementRoadSafetyColon848484849292929289898989656565657777777772.672.672.672.697.897.897.897.893.393.393.393.367.267.267.267.283.283.283.283.283.983.983.983.997.697.697.697.688.188.188.188.165.065.065.065.080.080.080.080.080.680.680.680.674.774.774.774.772.172.172.172.155.355.355.355.373.173.173.173.1%
Our methodXGBoostLiteratureRandom forestFormula size
Figure 1: The average test accuracies obtained with our method, random forests and XGBoost for all datasets. For all but the Colon dataset, also standard deviations are reported. For the UCI datasets we also include a comparison to the formula size method. When available, we have also included accuracies reported in literature, though these are not directly comparable due to different technical particularities in the experiments.
We compared our method empirically to random forests and XGBoost. These two comparison methods are very commonly used and state-of-the-art for tabular data. We tested all three methods on $12$ tabular datasets. The datasets can be categorized as follows.
@@
    $4$ tabular-data benchmarks out of $7$ binary classification benchmarks that were presented in the paper [10]. These datasets were also originally from the UCI machine learning repository.
@@
As mentioned in Section 3, our method utilizes readily available feature selection methods. We used Scikit-learn’s SelectKBest, which returns $k$ features that have the highest scores based on a given univariate statistical test. SelectKBest supports three methods for calculating the scores for the features: F-test, mutual information, and $\chi^{2}$-test. When generating formulas, we tested all three of these methods and selected the best in terms of validation accuracy. We emphasize that feature selection was performed after Booleanizing the data.
@@ 4.2 Results
Figure 1 contains the average test accuracies and standard deviations that were obtained by using our new method, random forests and XGBoost using ten-fold cross-validation. The left side of Figure 1 also contains results obtained with the formula-size method of [14]. We also report accuracies that we found in the literature for these datasets. For the left side of Figure 1, we include the accuracies that were reported for these datasets in the UCI machine learning repository. For the right side of Figure 1, for the case of Colon, we used the result reported in [23], while for remaining datasets we reported the accuries given in the full-version of [10].
@@
From Figure 1 we see that our method produces classifiers that mostly have accuracies similar to the ones obtained by the alternative methods. In particular, on medical data such as BreastCancer and Hepatitis, our method compares well to the reference methods. Perhaps the biggest gaps are found in the cases of CoverType and Electricity, two of the benchmark datasets. However, we emphasize that in [10], it was mentioned that all of the benchmark datasets were specifically selected to be “not too easy”. Hence, one should not perhaps expect that “simple” classifiers with focus on interpretability will perform too well with these datasets.
@@
We list the average runtime of our method on a single split of the data in the Appendix 6.3. For six of the datasets, the average runtime was less than 30 seconds. Of the remaining datasets, four had runtimes of less than 12 minutes. Even for the two largest datasets, Covertype and RoadSafety, the runtime is still less than an hour and, notably, systematically less than that of the hyperparameter optimizations of random forests and XGBoost. The experiments on our method, random forests and XBGoost were run on a laptop.
@@
We turn our attention to our main goal of interpretability. In Table 1 we report the average of the smallest number of features sufficient to reach the reported accuracy. Table 1 also gives the number of rows and attributes in the Booleanized datasets as well as the original datasets. We see from Table 1 that for eight out of 12 datasets, the best accuracy was reached already with on average less than four features, leading to very short formulas. Such short DNF-formulas are globally interpretable as follows. To interpret a short DNF-formula, one looks at each of the few conjunctions in the formula individually to determine their meaning. This is easy as each conjunction has very few attributes. The meaning of the entire formula is then given by the fact that it accepts only the cases given by the conjunctions and rejects everything else.
@@
Next we discuss the results obtained using the formula-size method of [14]. By the size of a propositional formula they mean the number of proposition symbols, conjunctions, disjunctions and negations in the formula. The results are given in Figure 1.
@@
We see that for most datasets on the left side in Figure 1, the accuracies obtained by the formula-size method are similar to our method. A meaningful difference can be seen in the HeartDisease dataset, where both the accuracy of 73.6 percent and standard deviation of 12.2 percent are significantly behind other methods. The formulas obtained are short, with formula size mostly less than 10.
@@
The downside of the formula-size method is computational resources. As described in the previous subsection, we used a Linux cluster featuring Intel Xeon 2.40 GHz CPUs with 8 CPU cores per run, employing a timeout of 72 hours per instance and a memory limit of 64 GB. For all but one dataset we ran the formula-size method on, the computation took more than 24 hours per split of the data with this high efficiency setup. Even for the easiest one, BreastCancer, the computation time was in the order of 12 hours per split. Furthermore, for the benchmark datasets and the high-dimensional biological dataset Colon, the formula-size method is unfeasible already for very small formula sizes and thus we do not report any results for those datasets.
@@
Data set Selected features Boolean features Original features Data points
@@ 4.3 Examples of obtained formulas
In the Appendix 6.4, there are further examples of formulas obtained as outputs of our method. For each dataset, we list the formula obtained for the first 90/10 split of our ten-fold cross-validation. While some of these formulas are quite long, they should still be rather fast to use for local explanations explicating why a particular input was classified in a particular way (cf. the Introduction for details on local explanations). For these experiments, we used the same maximum number 10 of features and the same limit of 1 percentage point to choose an accurate enough formula. These are hyperparameters that one could optimize to obtain a better balance of interpretability and accuracy on any specific dataset. For some datasets, raising the number of maximum features could lead to more accurate, but longer, formulas. Raising the percentage threshold on the other hand would lead to shorter, more interpretable formulas at the cost of some accuracy.
@@ 5 Conclusions
Our approach need not limit to Boolean formulas only, as we can naturally extend our work to general relational data. We can use, e.g., description logics and compute concepts $C_{1},\dots,C_{k}$ and then perform our procedure using $C_{1},\dots,C_{k}$, finding short Boolean combinations of concepts. This of course differs from the approach of computing empirical ideal classifiers in the original description logic, but can nevertheless be fruitful and interesting. We leave this for future work.